# Efficient and Intuitive Two-Phase Validation Across Multiple Models via Principal Components

### v1

@@ 4.1 About the Survey and Data
The NHANES includes various demographic, examination, laboratory, and questionnaire data. We focus on the dietary interview portion, called the “What We Eat in America” (WWEIA), which is conducted by the National Center for Health Statistics in collaboration with the U.S. Department of Agriculture and the U.S. Department of Health and Human Services. WWEIA captures individual-level daily consumption of $64$ nutrients and food components and is an incredibly valuable source for dietary intake data [Hoy2015].
In the most recent WWEIA (2021202120232023), a representative sample of individuals of all ages was asked to complete two $24$-hour dietary intake recalls approximately $1$ week apart. The food components and quantities recorded were converted into dietary nutrients using the Food and Nutrient Database for Dietary Studies, which provides caloric and nutrient information [fndds]. The NHANES data set reports the cumulative intake of these nutrients for each recall.
@@ 4.2 Modeling Diet-Driven Health Outcomes
We considered five models of interest, each one linking an examination or laboratory outcome $Y_{j}$ to a corresponding dietary intake exposure $X_{j}$. Each model was adjusted by the individuals’ sex, age, race and ethnicity (Hispanic origin), and education level to account for possible confounding, denoted by $\mathbf{Z}$}. Each model followed the linear regression in (1).
@@
$Y_{1}:$ Vitamin D $X_{1}:$ Calcium Vitamin D facilitates calcium absorption [CHRISTAKOS201125].
$Y_{2}:$ Resting Heart Rate $X_{2}:$ Caffeine Caffeine can affect heart rate and cardiovascular function [TURNBULL2017165].
$Y_{3}:$ High-Density Lipoproteins (HDL) Cholesterol $X_{3}$: Saturated Fat Saturated fat intake increases HDL cholesterol [Brinton1990].
$Y_{4}:$ Insulin $X_{4}$: Alcohol Alcohol consumption may influence insulin sensitivity [Schrieks2015].
$Y_{5}:$ Folate $X_{5}:$ Food Folate Low folate levels can lead to higher odds of anemia [Morris2007].
Table 1: The outcome of interest, associated dietary nutrient value, and the associated source data for the five models from the NHANES data. All covariates are sourced from the dietary interview portion of the National Health and Nutrition Examination Survey (NHANES), called “What We Eat in America.”
We focused on individuals who (i) completed the first telephone interview and (ii) were not missing any of the models’ outcomes, exposures, or covariates. A sample of $N=2388$ met these criteria and were used for analysis. While excluding individuals with missing data is not always advisable (e.g., when missing data patterns may be informative), we elected to do so here to simplify our illustration of the proposed validation study design.
@@ 4.3 Measurement Error in Dietary Intake
The NHANES dietary interview uses the United States Department of Agriculture Automated Multiple-Pass Method (AMPM) which is considered the most sophisticated methodology for 24-hour food recall [THOMPSON20175]. ADPM begins with a list of all foods and beverages consumed, followed by probing questions, collecting the time and occasion of meals, detailed descriptions, and a final nudge for anything else consumed.
@@
Two main limitations of the AMPM exist. First, it requires multiple days of recalls to obtain accurate mean measurements of intake, which is both costly and time-consuming. Second, it can induce self-report bias from psycho-social behaviors that result in under or over estimation of nutrient and energy intake, though underreporting is most common (see [THOMPSON20175] and references therein). Weighed food records are the gold standard for recording portions and the most accurate measurements, but has a high burden on participants. [willett2012nutritional]. Weighed food records require measuring all portions prior to consumption and any food waste that remains after a meal. A large-scale study such as the NHANES cannot implement weighed food records due to the size, cost, and burden. It is possible that measurement error may occur during the food recall and may be considered to impact assessment on the relationship of dietary intake with other biomarkers.
@@
We treat the available dietary intake exposures from the NHANES dataset as the error-free versions $\boldsymbol{X}=(X_{1},X_{2},X_{3},X_{4},X_{5})$. Calcium ($X_{1}$) and saturated fat ($X_{3}$) were moderately correlated with each other ($\rho=0.59$) and weakly correlated with food folate ($X_{5}$) ($\rho=0.38$ and $0.34$, respectively). Caffeine ($X_{2}$) and alcohol ($X_{4}$) exhibited minimal correlation with other exposures (all $\rho\leq 0.16$). The Web Appendix contains the full correlation matrix.
@@
We simulated the error-prone dietary intake exposures $\boldsymbol{X}^{*}=(X_{1}^{*},X_{2}^{*},X_{3}^{*},X_{4}^{*},X_{5}^{*})$ by generating the measurement errors $\boldsymbol{U}=(U_{1},U_{2},U_{3},U_{4},U_{5})$ from a multivariate normal distribution with mean vector $\mathbf{0}_{5}$ and covariance matrix $\mathrm{diag}(77805.2,7630.9,70.5,119.1,4991.9)$, where $\mathrm{diag}(\cdot)$ denotes a diagonal matrix. For each $X_{j}$, the variance of the error $U_{j}$ was chosen to be $\sigma_{uj}^{2}=\widehat{\textrm{Var}}(X_{j})/4$, i.e., $1/4$ times the estimated variance of the corresponding error-free exposure. Then, we combined the error-free versions from NHANES and simulated errors to get $\boldsymbol{X}^{*}=\boldsymbol{X}+\boldsymbol{U}$. Scatterplots of $\boldsymbol{X}$ versus $\boldsymbol{X}^{*}$ are in the Web Appendix.
@@
As anticipated, after adding measurement error, the correlations between the five dietary intake covariates all exposures (Web Appendix). For example, the correlation between $X_{1}^{*}$ (calcium) and $X_{3}^{*}$ (saturated fat) was $\rho^{*}=0.46$, while that between their error-free versions was $\rho=0.59$. We expected that the proposed ETS-$PC_{1}^{*}$ design would be beneficial in these data due to the relationships between these exposures.
@@ 4.4 Partial Validation of Dietary Intake Covariates
For our purposes, all $N=2388$ individuals in the NHANES dataset had outcomes $\boldsymbol{Y}=(Y_{1},Y_{2},Y_{3},Y_{4},Y_{5})$, error-prone dietary intake $\boldsymbol{X}^{*}$, and other covariates $\boldsymbol{Z}$ observed. However, only a subset of $n=250$ of them underwent “validation” and had error-free dietary intake $\boldsymbol{X}$ observed. We considered three strategies to select this subset of validated patients.
First, we included the naive SRS design, where the $250$ individuals were chosen for validation from the original sample of $2388$ with equal probability (i.e., ignoring all information). While theoretically valid, this design leads to the widest 95% confidence intervals (95% CIs) for ${\boldsymbol{\beta}}_{1}$. It leads to model estimates that are asymptotically consistent (i.e., converge in probability to the true coefficients) as both sample sizes ($N$ and $n$) go to infinity. However, there can be discrepancies between the SRS estimates and full validation in finite samples.
Second, we considered an ETS-$X_{1}^{*}$ design that focuses on the error-prone dietary intake exposure, $X_{1}^{*}$ (calcium), from the first model. This approach might appeal if we were primarily interested in this model, so that we tailored our validation sampling strategy to focus on it. However, we may later decide that the other four models are also relevant, and then re-purpose the validation study data to estimate them. The ETS-$X_{1}^{*}$ design should offer the narrowest 95% CI for Model 1, but it may not even outperform SRS for Models 2–5.
Third, we applied our proposed ETS-$PC_{1}^{*}$ design, which prioritized individuals with the most extreme values of the first principal component fit to the five error-prone exposures. The first principal component, $PC_{1}^{*}$, explained $35\%$ of the variability of the original five variables $\boldsymbol{X}^{*}$. All error-prone nutrient intake exposures contributed to $PC_{1}^{*}$, with the largest loading of $0.59$ for $X_{1}^{*}$ (calcium) and the smallest of $0.14$ for $X_{4}^{*}$ (alcohol).
For comparison, we calculated the average percent variability explained by $PC_{1}^{*}$ across replications in our simulations. When varying the covariance structures for $\boldsymbol{X}$, $28\%$ was explained under independence, $48\%$ under equal dependence, and $39\%$ under unequal dependence. As the exposure error variance $\sigma_{U}^{2}$ increased, the percent of variability explained by $PC_{1}^{*}$ dropped from $64\%$ to $48\%$. Based on the correlation matrix for the NHANES data (Web Appendix), their covariance structure resembled the unequal dependence scenario; the $35\%$ variability explained by $PC_{1}^{*}$ aligned well with the simulations, too. Plus, the magnitude of measurement errors in the nutrient intake exposures were fairly severe (Web Appendix).
@@ 4.5 Estimates from Multiple Models
In Figure 3A, we summarize estimates from the five diet-driven health outcome models from Section 4.2. First, we computed the “gold standard” model estimates using the error-free covariates $X_{1},\dots,X_{5}$ on all $N=2388$ individuals from the original NHANES data. The estimated slopes $\widehat{\beta}_{1j}$ for each of the five nutrient intake exposures was approximately 0.
Figure 3: A) Coefficient estimates and corresponding $95\%$ confidence intervals and B) widths of $95\%$ confidence intervals for $\beta_{1j}$, corresponding to each of the five nutrient intake exposures ($j\in\{1,\dots,5\}$) from the five diet-driven health outcome models. In addition to the gold standard analysis (with all individuals having validation data), we considered partial validation with simple random sampling (SRS), extreme tail sampling on $X_{1}^{*}$ (ETS-$X_{1}^{*}$), and extreme tail sampling on $PC_{1}^{*}$ (ETS-$PC_{1}^{*}$). The analyses with partial validation data used multiple imputation.
Then, we considered analyses with a validation study of $n=250$ individuals, sampled following the three designs in Section 4.4. For the remaining $2138$ unvalidated individuals, error-free nutrient intake exposures were multiply imputed $75$ times, with the imputation models tailored to the different sampling strategies. In general, the three imputation analyses led to estimates for $\beta_{1j}$ that differed slightly from the gold standard values. However, they would become more similar if more validation data were available (i.e., $n$ increased).
The relative widths of the 95% CIs are provided in Figure 3B. The ETS-$PC_{1}^{*}$ design offered the narrowest 95% CIs in all five models. It had the largest advantages over SRS and ETS-$X_{1}^{*}$ in Model $4$, which was interesting given that the fourth nutrient intake exposure $X_{4}^{*}$ (alcohol) was not very correlated with any of the others and had the smallest loading on $PC_{1}^{*}$. The ETS-$X_{1}^{*}$ design did not offer improvement over the SRS design for the first model (or the fourth) but led to slightly narrower 95% CIs for the others. It was particularly surprising for ETS-$X_{1}^{*}$ to not perform best for the corresponding model with $X_{1}$, as it did in simulations. However, the error-prone nutrient intake exposure could be so noisy that sampling on the extremes of $X_{1}^{*}$ did not necessarily lead us to validate the extremes of $X_{1}$.
@@ 5 Discussion
Extreme tail sampling on the first principal component offers an accessible, intuitive way to balance multiple, competing analytical priorities when choosing which individuals to validate in studies with exposure measurement error. This strategy amplifies the benefits of a conventional single-objective two-phase design by avoiding the expenses of conducting separate, targeted validation samples for distinct models. It does so without incurring computational complexity, and we provide a user-friendly R package.
In simulations, our proposed sampling strategy improved the overall efficiency of the two-phase estimates across multiple models of interest under mild assumptions. The improvement was evident under varying covariance structures, measurement error severities, and validation proportions. Models where the error-prone exposure was strongly correlated with the others saw the greatest improvement. Still, when the exposures’ variability was relatively small compared to the outcomes’, any ETS procedure (on a single exposure or principal component) will see decreased benefits as the improved precision of covariates will not substantially improve the precision of the overall effect estimates.
We demonstrate the design’s advantages in multiple imputation analyses, where model specification is critical to achieving unbiased estimates. In practice, selection procedures (e.g., based on information criteria) can help select the “best-fitting” imputation model [SCHOMAKER2014758]. However, the advantages of sampling on the first principal component should also manifest if applied to other model-based estimation procedures [lotspeich2023optimal], which opens up possibilities for more robust analytical approaches, like semiparametric MLE [tao2021].
Toward broader applicability of the proposed sampling strategy, there are multiple interesting directions for future work. First, investigations into efficiency gains with other common outcome types (including binary, count, or time-to-event) are needed. Second, adopting existing informative or even optimal validation study designs for a dimension-reduced version of the error-prone exposures (like the first principal component) could offer further advantages. Third, modifications to other error settings (e.g., error-prone outcomes) are promising; in particular, the set of variables contributing to the PCA will need to be revisited. Finally, approaches considering more than the first principal component (e.g., splitting the validation study into $k$ further subsamples, and choosing $n/k$ patients from the extremes of the first through $k$th principal component).

### v2

@@ 4.1 About the Survey and Data
The NHANES includes various demographic, examination, laboratory, and questionnaire data. We focus on the dietary interview portion, called the “What We Eat in America” (WWEIA), which captures individual-level daily consumption of $64$ nutrients and food components and is an incredibly valuable source for dietary intake data [16].
In the most recent WWEIA (2021202120232023), a representative sample of individuals of all ages was asked to complete two $24$-hour dietary intake recalls approximately $1$ week apart. The food components and quantities recorded were converted into dietary nutrients using the Food and Nutrient Database for Dietary Studies, which provides caloric and nutrient information [40]. The NHANES data set reports the cumulative intake of these nutrients for each recall.
@@ 4.2 Modeling Diet-Driven Health Outcomes
We considered five models of interest, each one linking an examination or laboratory outcome $Y_{j}$ to a corresponding dietary intake exposure $X_{j}$ (Table 1). Each model was adjusted by the individuals’ sex, age, race and ethnicity (Hispanic origin), and education level to account for possible confounding, denoted by $\mathbf{Z}$}. Each model followed the linear regression in (1), and estimating the exposure coefficients ${\boldsymbol{\beta}}_{1}$ across all models was our goal.
@@
$Y_{1}:$ Vitamin D $X_{1}:$ Calcium Vitamin D facilitates calcium absorption [5].
$Y_{2}:$ Resting Heart Rate $X_{2}:$ Caffeine Caffeine can affect heart rate and cardiovascular function [39].
$Y_{3}:$ High-Density Lipoproteins (HDL) Cholesterol $X_{3}$: Saturated Fat Saturated fat intake increases HDL cholesterol [4].
$Y_{4}:$ Insulin $X_{4}$: Alcohol Alcohol consumption may influence insulin sensitivity [33].
$Y_{5}:$ Folate $X_{5}:$ Food Folate Low folate levels can lead to higher odds of anemia [27].
Table 1: The outcome of interest, associated dietary nutrient value, and the associated source data for the five models from the NHANES data. All exposures are sourced from the dietary interview portion of the National Health and Nutrition Examination Survey (NHANES), called “What We Eat in America.”
We focused on individuals who (i) completed the first telephone interview and (ii) were not missing any of the models’ outcomes, exposures, or confounders. A sample of $N=2388$ met these criteria and were used for analysis. While excluding individuals with missing data is not always advisable (e.g., when missing data patterns may be informative), we elected to do so here to simplify our illustration of the proposed validation study design.
@@ 4.3 Measurement Error in Dietary Intake
The NHANES dietary interview uses the Automated Multiple-Pass Method (AMPM) which is considered the most sophisticated methodology for 24-hour food recall [38]. ADPM begins with a list of all foods and beverages consumed, followed by probing questions, collecting the time and occasion of meals, detailed descriptions, and a final nudge for anything else consumed.
@@
Two main limitations of the AMPM exist. First, it requires multiple days of recalls to obtain accurate mean measurements of intake, which is both costly and time-consuming. Second, it can induce self-report bias from psycho-social behaviors that result in under or over estimation of nutrient and energy intake, though underreporting is most common (see [38] and references therein). Weighed food records are the gold standard for recording portions and the most accurate measurements, but has a high burden on participants. [44]. Weighed food records require measuring all portions prior to consumption and any food waste that remains after a meal. A large-scale study such as the NHANES cannot implement weighed food records due to the size, cost, and burden. Measurement error may occur during the food recall and impact assessment on the relationship of dietary intake with other biomarkers.
@@
We treat the available dietary intake exposures from the NHANES dataset as the error-free versions $\boldsymbol{X}=(X_{1},X_{2},X_{3},X_{4},X_{5})$. Calcium ($X_{1}$) and saturated fat ($X_{3}$) were moderately correlated with each other ($\hat{\rho}=0.59$) and weakly correlated with food folate ($X_{5}$) ($\hat{\rho}=0.38$ and $0.34$, respectively). Caffeine ($X_{2}$) and alcohol ($X_{4}$) exhibited minimal correlation with other exposures (all $\hat{\rho}\leq 0.16$). See Supplemental Figure S5A for the full correlation matrix.
@@
We simulated error-prone $\boldsymbol{X}^{*}=(X_{1}^{*},X_{2}^{*},X_{3}^{*},X_{4}^{*},X_{5}^{*})$ by generating measurement errors $\boldsymbol{U}=(U_{1},U_{2},U_{3},U_{4},U_{5})$ from a multivariate normal distribution with mean vector $\mathbf{0}_{5}$ and covariance matrix $\mathrm{diag}(77805.2,7630.9,70.5,119.1,4991.9)$, where $\mathrm{diag}(\cdot)$ denotes a diagonal matrix. The error variance was chosen to be $\sigma_{Uj}^{2}=\widehat{\textrm{Var}}(X_{j})/4$, i.e., $1/4$ times the estimated variance of the corresponding error-free exposure. Then, we combined added these errors to the NHANES data to get $\boldsymbol{X}^{*}=\boldsymbol{X}+\boldsymbol{U}$. Supplemental Figure S6 contains scatterplots of $\boldsymbol{X}$ versus $\boldsymbol{X}^{*}$.
@@
As anticipated, after adding measurement error, the correlations between the five dietary intake covariates all exposures decreased but persisted (Supplemental Figure S5B). For example, the correlation between $X_{1}^{*}$ (calcium) and $X_{3}^{*}$ (saturated fat) was $\hat{\rho}^{*}=0.46$, while that between their error-free versions was $\hat{\rho}=0.59$.
@@ 4.4 Partial Validation of Dietary Intake Covariates
All $N=2388$ individuals had outcomes $\boldsymbol{Y}=(Y_{1},Y_{2},Y_{3},Y_{4},Y_{5})$, error-prone dietary intake $\boldsymbol{X}^{*}$, and other confounders $\boldsymbol{Z}$ observed. However, only a subset of $n=250$ of them underwent “validation” and had error-free dietary intake $\boldsymbol{X}$ observed.
First, we included the SRS design, which, while theoretically valid, leads to the widest 95% confidence intervals ($95\%$ CIs) for ${\boldsymbol{\beta}}_{1}$. It leads to model estimates that are asymptotically consistent (i.e., converge in probability to the true coefficients) as both sample sizes ($N$ and $n$) go to infinity. However, there can be discrepancies between the SRS estimates and full validation in finite samples.
Second, we considered an ETS-$X_{1}^{*}$ design that focuses on the error-prone dietary intake exposure, $X_{1}^{*}$ (calcium). This approach might appeal if we were primarily interested in this model, so that we tailored our validation sampling strategy to focus on it. However, we may later decide that the other four models are also relevant, and then re-purpose the validation study data to estimate them. The ETS-$X_{1}^{*}$ design should offer the narrowest $95\%$ CI for Model 1, but it may not even outperform SRS for Models 2–5.
Third, we applied our proposed ETS-$PC_{1}^{*}$ design. The first principal component, $PC_{1}^{*}$, explained $35\%$ of the variability of $\boldsymbol{X}^{*}$. All error-prone nutrient intake exposures contributed to $PC_{1}^{*}$, with the largest loading of $0.59$ for $X_{1}^{*}$ (calcium) and the smallest of $0.14$ for $X_{4}^{*}$ (alcohol). For comparison, we calculated the average percent variability explained by $PC_{1}^{*}$ in simulations. When varying the covariance structures for $\boldsymbol{X}$, $28\%$ was explained under independence, $48\%$ under equal dependence, and $39\%$ under unequal dependence. As the exposure error variance $\sigma_{U}^{2}$ increased, the percent of variability explained by $PC_{1}^{*}$ dropped from $64\%$ to $48\%$. Based on the correlation matrix for the NHANES data (Supplemental Figure S5), their covariance structure resembled the unequal dependence scenario; the $35\%$ variability explained by $PC_{1}^{*}$ aligned well with the simulations, too. Plus, the magnitude of measurement errors in the nutrient intake exposures was fairly severe.
@@ 4.5 Estimates from Multiple Models
Figure 6A summarizes estimates from the five diet-driven health outcome models. First, we computed the “gold standard” model estimates using the error-free covariates $\boldsymbol{X}$ on all $N=2388$ individuals from the original NHANES data. The estimated slopes $\widehat{\beta}_{1j}$ for each of the five nutrient intake exposures was approximately 0. All other coefficient estimates are summarized in Table 2.
Figure 6: Coefficient estimates and corresponding $95\%$ confidence intervals (A) and relative widths of $95\%$ confidence intervals for $\beta_{1j}$, corresponding to each of the five nutrient intake exposures ($j\in\{1,\dots,5\}$) from the five diet-driven health outcome models (B). In addition to the gold standard analysis (with all individuals having validation data), we considered partial validation with simple random sampling (SRS), extreme tail sampling on $X_{1}^{*}$ (ETS-$X_{1}^{*}$), and extreme tail sampling on $PC_{1}^{*}$ (ETS-$PC_{1}^{*}$). The analyses with partial validation data used multiple imputation.
Then, we considered analyses with a validation study of $n=250$ individuals. For the remaining $2138$ unvalidated individuals, error-free nutrient intake exposures were multiply imputed $75$ times. In general, the three imputation analyses led to estimates for $\beta_{1j}$ that differed slightly from the gold standard values. However, they would become more similar if more validation data were available.
The total coefficient variability for the gold standard analysis was $\sum_{j=1}^{5}\widehat{V}(\hat{\beta}_{1j})=0.002$, while for the partially validated analyses using the SRS and ETS-$X_{1}^{*}$ designs it was =0.007=0.007. With the ETS-$PC_{1}^{*}$ design, the total variability of $0.004$ with the multiple imputation analysis was much closer to the gold standard. Additionally, the ETS-$PC_{1}^{*}$ design offered narrower $95\%$ CIs in all five models than the SRS or ETS-$X_{1}^{*}$ (Figure 6B). It had the largest advantages in Model $4$, which was interesting given that the fourth nutrient intake exposure $X_{4}^{*}$ (alcohol) was not very correlated and had the smallest loading on $PC_{1}^{*}$. The ETS-$X_{1}^{*}$ design did not offer much improvement over SRS. It was particularly surprising for ETS-$X_{1}^{*}$ to not perform best for the corresponding model with $X_{1}$, as it did in simulations. However, the error-prone nutrient intake exposure could be so noisy that sampling on the extremes of $X_{1}^{*}$ did not necessarily lead us to validate the extremes of $X_{1}$ (which are the source of the efficiency gain).
@@ 5 Discussion
In simulations, our proposed sampling strategy improved the overall efficiency of the two-phase estimates across multiple models of interest under mild assumptions. The improvement was evident under varying covariance structures, measurement error severities, and validation proportions. Models where the error-prone exposure was strongly correlated with the others saw the greatest improvement. Still, when the exposures’ variability was relatively small compared to the outcomes’, any ETS procedure will see decreased benefits as the improved precision of exposures will not substantially improve the precision of the overall effect estimates.
We demonstrate the ETS-$PC_{1}^{*}$ design’s advantages in single and multiple imputation analyses, where model specification is critical to achieving unbiased estimates. In practice, selection procedures (e.g., based on information criteria) can help select the “best-fitting” imputation model [32]. However, the advantages of sampling on the first principal component should also manifest if applied to other model-based estimation procedures [20], which opens up possibilities for more robust analytical approaches, like semiparametric MLE [36].
Toward broader applicability of the proposed sampling strategy, there are multiple interesting directions for future work. First, investigations into efficiency gains with other common outcome types (including binary, count, or time-to-event) are needed. Second, adopting other existing designs for a dimension-reduced version of the error-prone exposures (like the first principal component) could offer further advantages. Third, modifications to other error settings (e.g., error-prone outcomes) are promising; in particular, the set of variables contributing to the PCA will need to be revisited.
6 Conclusion
Extreme tail sampling on the first principal component offers an accessible, intuitive way to balance multiple, competing analytical priorities when choosing which individuals to validate in studies with exposure measurement error. This strategy amplifies the benefits of a conventional single-objective two-phase design by avoiding the expenses of conducting separate, targeted validation samples for distinct models. It does so without incurring computational complexity, and we provide the user-friendly R auditDesignR package at https://github.com/sarahlotspeich/auditDesignR.